# Highest weight category structures on Rep(B) and full exceptional collections on generalized flag varieties over $\mathbb{Z}$

### v2

@@ 13.2 The set $\{\hat{M}_{v}\}_{v\in W^{P}}$
in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))$. Take $\hat{M}_{v}=Q(e_{v})$ for $p\prec v\in W^{P}$ and $\hat{M}_{v}=P(-e_{v})^{*}$ for $p\succ v\in W^{P}$. It does not matter what we choose for $\hat{M}_{p}$ itself.
Using Theorem 13.11 instead of Theorem 10.1, we find that ${\mathsf{cone}}(\hat{f}_{p})$ lies in the triangulated hull of $\{\hat{M}_{v}\otimes\nabla_{\lambda}\mid v\in W^{P},\ e_{v}<_{a}e_{p},\ \lambda\in X(T)_{+}\}$. So ${\mathsf{cone}}(\hat{f}_{p})$ lies in the triangulated hull of $\hat{\sf P}_{\prec p}\cup\hat{\sf Q}_{\succ p}$. One uses this to show that $\hat{X}_{p}$ and $\hat{Y}_{p}$ are isomorphic. Mutatis mutandis the old constructions and proofs go through and Theorems 13.2, 13.3, 13.4 follow.
@@ 13.3 A pairing
Let $X=G/B$, $Y=G/P$ and let $\pi:X\to Y$ be the projection. It induces a map
$\pi^{*}:K_{T}(Y)\to K_{T}(X)$ and therefore, by subsection 3.3, a pairing
$K_{T}(Y)\times K_{T}(X)\to R(T),\quad([{\mathcal{F}}],[{\mathcal{G}}])\mapsto\langle\pi^{*}[{\mathcal{F}}],[{\mathcal{G}}]\rangle.$
For $v,w\in W^{P}$ with $vw_{0}\not\leq ww_{0}$ the pairing sends $([\mathcal{L}_{Y}(P(-e_{v}))],[\mathcal{O}_{X^{ww_{0}}}(-\partial X^{ww_{0}})])$ to $\beta_{v,ww_{0}}=0$. And it sends $([\mathcal{O}_{\overline{B^{+}vP}/P}],[\mathcal{O}_{X^{ww_{0}}}(-\partial X^{ww_{0}})])$ to $\delta_{vw}$ for $v,w\in W^{P}$. If ${\sf k}={\mathbb{C}}$ it follows that $[\mathcal{L}_{Y}(P(-e_{v}))]$ is in the $R(T)$-span of the Schubert classes $[\mathcal{O}_{\overline{B^{+}wP}/P}]$ with $w\in W^{P}$, $w\leq v$.
@@ 14.1 Borel-Weil-Bott theorem
    If $\lambda\in\overline{C}_{\mathbb{Z}}$ with $\lambda\notin X({T})_{+}$, then
    ${\rm H}^{i}({G}/{B},\mathcal{L}_{w\cdot\lambda})=0$
    for all $w\in W$. Here $\cdot$ is the dot–action of the Weyl group $W$ on $X({T})$.
@@
    If $\lambda\in\overline{C}_{\mathbb{Z}}\cap X({T})_{+}$, then for all $w\in W$ and $i\in\mathbb{N}$
    ${\rm H}^{i}({G}/{B},\mathcal{L}_{w\cdot\lambda})={\rm H}^{0}({G}/{B},\mathcal{L}_{\lambda}),$
    if $i=l(w)$ and otherwise ${\rm H}^{i}({G}/{B},\mathcal{L}_{w\cdot\lambda})=0$. Here $l(w)$ is the length function.
@@ 14.2 Kempf vanishing theorem
[29, Part II, Proposition 4.5] and [18, Section 2.6] If $\lambda\in X({T})_{+}$, then
${\rm H}^{i}({G}/{B},\mathcal{L}_{\lambda})=0$
@@
By [3, Corollary 2.6, (ii)], ${\rm H}^{0}({\mathbb{G}}/{\mathbb{B}},{\mathbb{L}}_{\lambda})$ are free $\mathbb{Z}$-modules for $\lambda\in X({\mathbb{T}})_{+}$. Combined with the universal coefficient theorem [29, II, Section 8.7, (1)] and Theorem 14.2, this gives ${\rm H}^{i}({\mathbb{G}}/{\mathbb{B}},{\mathbb{L}}_{\lambda})=0$ for $i>0$.
@@ 14.3 Kodaira vanishing theorem
[7, Theorem 1.2.9] Let $k$ be a field of characteristic $p>0$ and $X$ be a Frobenius split $k$-variety. Assume given an ample line bundle $\mathcal{L}$ on $X$. Then $H^{i}(X,\mathcal{L}\otimes\omega_{X})=0$ for $i>0$.
Generalized flag varieties $G/P$ are Frobenius split, [7, Theorem 2.2.5], thus Kodaira vanishing holds for these.
@@ 14.4 Demazure functors
By [29, Section 5.17-5.18], the Joseph-Demazure functors ${\sf H}_{w}$ from Section 2.2 can be related to functors which are defined geometrically. Let $R$ be the root system of a simple simply connected group $G$ of rank $r$ and $\Pi=\{\alpha_{1},\dots,\alpha_{r}\}\subset R$ the simple roots. Let $s_{\alpha_{i}},i=1,\dots,r$ be the associated simple reflections of the Weyl group $W$. Given a simple root $\alpha$, let $\pi_{\alpha}:{G}/{B}\rightarrow{G}/{P}_{\alpha}$ denote the projection. Associated to $\alpha$ is an endofunctor ${\rm R}^{0}{\sf D}_{\alpha}:=\pi_{\alpha}^{\ast}{\rm R}^{0}{\pi_{\alpha}}_{\ast}$ of ${\mathsf{QCoh}}^{G}({G}/{B})$ (the functor $\pi_{\alpha}^{\ast}$ is exact, the morphism $\pi_{\alpha}$ being smooth). Let $w$ be an element of the Weyl group. For a reduced expression of $w=s_{\alpha_{1}}s_{\alpha_{2}}\dots s_{\alpha_{n}}$ consider the functor ${\rm R}^{0}{\sf D}_{w}={\rm R}^{0}{\sf D}_{\alpha_{1}}\circ\dots\circ{\rm R}^{0}{\sf D}_{\alpha_{n}}=\pi_{\alpha_{1}}^{\ast}{\rm R}^{0}{\pi_{\alpha_{1}}}_{\ast}\dots\pi_{\alpha_{n}}^{\ast}{\rm R}^{0}{\pi_{\alpha_{n}}}_{\ast}$. The functor ${\rm R}^{0}{\sf D}_{w}$ is also a left exact functor.
Similarly to the case of ${\sf H}_{w}$, let ${\sf D}_{w}={\sf D}_{\alpha_{1}}\circ\dots\circ{\sf D}_{\alpha_{n}}=\pi_{\alpha_{1}}^{\ast}{\pi_{\alpha_{1}}}_{\ast}\dots\pi_{\alpha_{n}}^{\ast}{\pi_{\alpha_{n}}}_{\ast}$ denote the right derived functor of ${\rm R}^{0}{\sf D}_{w}$; it is an endofunctor of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}({\mathsf{QCoh}}^{G}({G}/{B}))$ [^4] Note that the functor ${\sf D}_{w}$ is also defined on the derived category of non-equivariant coherent sheaves on $G/B$ and eventually will be considered on that category.. Since the morphisms $\pi_{\alpha_{i}}$’s are smooth and proper, the functors ${\sf D}_{w}$’s restrict to endofunctors of the bounded derived category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\sf Coh}^{G}({G}/{B}))$ [^5] The same footnote as above applies..
Now set ${G}:={B},X={G}$ (res., ${G}:={P}_{\alpha},X={G}$) in the definition of the sheafification functor $\mathcal{L}$ and denote $\mathcal{L}_{B}$ (resp., $\mathcal{L}_{{P}_{\alpha}}$) the corresponding functors ${{\mathrm{Rep}}}({B})\rightarrow{\mathsf{QCoh}}^{G}({G}/{B})$ (resp., ${{\mathrm{Rep}}}({P}_{\alpha})\rightarrow{\mathsf{QCoh}}^{G}({G}/{P}_{\alpha}$)).
@@
Let $M\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}({{\mathrm{Rep}}}(B))$ (resp., $M\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({{\mathrm{rep}}}(B))$). Then there is an isomorphism $\mathcal{L}_{B}({\rm R}{\sf H}_{w}(M))={\sf D}_{w}\mathcal{L}_{B}(M)$.
@@
See [35, Lemma 5]. ∎
To unburden the notation, let $X$ denote ${G}/{B}$. Occasionally though when there is a need, we will stress the notation ${G}/{B}$. For a simple root $\alpha_{i}$ let ${P}_{\alpha_{i}}$ denote the minimal parabolic generated by ${B}$ and $U_{\alpha_{i}}$, and denote $Y_{i}:={G}/{P}_{\alpha_{i}}$ with $\pi_{\alpha_{i}}:X\rightarrow Y_{i}$. Given a simple root $\alpha_{i}$, we denote $\sf D_{i}$ the Demazure operator associated to $\alpha_{i}$, i.e. $\sf D_{i}:=\sf D_{\alpha_{i}}$. Throughout this section, we employ the notation of Section 1.5. The following theorem is the starting point for unraveling the category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({{\mathrm{rep}}}({B}))$ as a module category over $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({{\mathrm{rep}}}(G))$:

### v3

@@ 13.2 The set $\{\hat{M}_{v}\}_{v\in W^{\mathbf{P}}}$
in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{P}))$. Take $\hat{M}_{v}=Q(e_{v})$ for $p\prec v\in W^{\mathbf{P}}$ and $\hat{M}_{v}=P(-e_{v})^{*}$ for $p\succ v\in W^{\mathbf{P}}$. It does not matter what we choose for $\hat{M}_{p}$ itself.
Using Theorem 13.12 instead of Theorem 10.1, we find that ${\mathsf{cone}}(\hat{f}_{p})$ lies in the triangulated hull of $\{\hat{M}_{v}\otimes\nabla_{\lambda}\mid v\in W^{\mathbf{P}},\ e_{v}<_{a}e_{p},\ \lambda\in X(\mathbf{T})_{+}\}$. So ${\mathsf{cone}}(\hat{f}_{p})$ lies in the triangulated hull of $\hat{\sf P}_{\prec p}\cup\hat{\sf Q}_{\succ p}$. One uses this to show that $\hat{X}_{p}$ and $\hat{Y}_{p}$ are isomorphic. Mutatis mutandis the old constructions and proofs go through and Theorems 13.2, 13.3, 13.5 follow. ∎
Remains to prove Theorem 13.7.
Lemma 13.17.
Let $s$ be a simple refection and $\mathbf{P}_{s}$ the corresponding minimal parabolic.
  • If $s\lambda<_{d}\lambda$ then $0\to Q(s\lambda)\to{{\sf H}_{s}}(Q(\lambda))\to Q(\lambda)\to 0$ is exact,
  • If $s\lambda=\lambda$ then ${{\sf H}_{s}}(Q(\lambda))=Q(\lambda)$,
  • If $s\lambda>_{d}\lambda$ then ${{\sf H}_{s}}(Q(\lambda))=0$.
Proof.
First assume $s\lambda\leq_{d}\lambda$. Then $\lambda$ is $\mathbf{P}_{s}$-dominant. The extremal weights of ${{\sf H}}_{s}(Q(\lambda))$ are the extremal weights $\lambda$, $s\lambda$ of ${{\sf H}}_{s}({\sf k}_{\lambda})$, because the weights $\nu$ of $Q(\lambda)$ different from $\lambda$ are too short to contribute, cf. proof of Lemma 13.10. By definition $Q(\lambda)$ is the kernel of a surjective map $P(\lambda)\rightarrow{\rm H}^{0}(\partial X_{w},\mathcal{L}(\lambda^{+}))$, and we get by Lemma 5.11 a relative Schubert filtration on ${{\sf H}}_{s}(Q(\lambda))$, with sections described by the extremal weights. Moreover, the submodule ${{\sf H}}_{s}({\sf k}_{\lambda})$ of ${{\sf H}}_{s}(Q(\lambda))$ has the $\mathbf{B}$-socle ${\sf k}_{s\lambda}$, so we know in which order the sections of the filtration occur. The first two parts follow.
Now assume $s\lambda>_{d}\lambda$. By [48, Theorem 1.9, (ii)], $Q(\lambda)$ is acyclic for ${{\sf H}}_{w},w\in W$. Applying ${{\sf H}}_{s}$ to the exact sequence $0\to Q(\lambda)\to{{\sf H}}_{s}(Q(s\lambda))\to Q(s\lambda)\to 0$, and using the idempotency ${{\sf H}}_{s}{{\sf H}}_{s}={{\sf H}}_{s}$ of ${{\sf H}}_{s}$’s, we conclude that ${{\sf H}}_{s}(Q(\lambda))=0$. ∎
Lemma 13.18.
Let $s$ be a simple refection and $\mathbf{P}_{s}$ the corresponding minimal parabolic. Let $w\in W$.
  • If $se_{w}<_{d}e_{w}$ then $ws\succ w$ and $0\to Q(e_{ws})\to{{\sf H}_{s}}(Q(e_{w}))\to Q(e_{w})\to 0$ is exact,
  • If $se_{w}=e_{w}$ then ${{\sf H}_{s}}(Q(e_{w}))=Q(e_{w})$,
  • If $se_{w}>_{d}e_{w}$ then ${{\sf H}_{s}}(Q(e_{w}))=0$.
Proof.
Recall that $w^{-1}$ is a minimal coset representative of the stabilizer in $W$ of the dominant weight $we_{w}$. If $se_{w}<_{d}e_{w}$, then $se_{w}>_{e}e_{w}$ by Lemma 8.21, so $\ell(sw^{-1})=\ell(w^{-1})+1$ and $sw^{-1}$ is also a minimal coset representative of the stabilizer in $W$ of the dominant weight $we_{w}$. So $sw^{-1}\alpha>0$ for simple roots $\alpha$ perpendicular to $we_{w}$. It follows that $se_{w}$ is the Steinberg weight $e_{ws}$. The other points are clear from Lemma 13.17. ∎
Lemma 13.19.
Let $w\in W^{\mathbf{P}}$. Then $\hat{Q}(e_{w})\to Q(e_{w})$ is surjective and its kernel lies in ${\sf Q}_{\succ p}$.
Proof.
Observe that $\hat{Q}(e_{w})$ is nonzero by Theorem 13.16. Recall there is a a reduced decomposition $w_{0}^{\mathbf{P}}=s_{1}\cdots s_{l}$ of $w_{0}^{\mathbf{P}}$ and that ${\mathrm{res}}^{\mathbf{P}}_{\mathbf{B}}{\mathrm{ind}}_{\mathbf{B}}^{\mathbf{P}}={\sf H}_{s_{1}}\circ\cdots\circ{\sf H}_{s_{l}}$. Therefore we may assume $\mathbf{P}$ is a minimal parabolic and apply Lemma 13.18, taking into account that all modules $Q(\lambda)$ are ${\sf H}_{s}$ acyclic for simple $s$. ∎
Lemma 13.20.
For $w\in W^{\mathbf{P}}$ one has $\hat{\sf Q}_{\succ p}\subset{\sf Q}_{\succ p}$.
Proof.
By descending induction along $\succ$, using Lemma 13.19. ∎
Proof of Theorem 13.7.
Take $p\in W^{\mathbf{P}}$. By equation (34) we have
$Y_{p}={i^{p}}_{\preceq p}^{\ !}(Q(e_{p})).$ (51)
By Lemma 13.19 we have
${i^{p}}_{\preceq p}^{\ !}(Q(e_{p}))={i^{p}}_{\preceq p}^{\ !}(\hat{Q}(e_{p})).$ (52)
As ${\sf cone}(\hat{Y}_{p}\rightarrow\hat{Q}(e_{p}))\in\hat{\sf P}_{\preceq p}^{\perp}=\hat{\sf Q}_{\succ p}\subset{\sf Q}_{\succ p}$, we also have
${i^{p}}_{\preceq p}^{\ !}(\hat{Q}(e_{p}))={i^{p}}_{\preceq p}^{\ !}(\hat{Y}_{p}).$ (53)
Further $\hat{Y}_{p}=\hat{X}_{p}\in\hat{\sf P}_{\preceq p}\subset{\sf P}_{\preceq p}$, so that
${i^{p}}_{\preceq p}^{\ !}(\hat{Y}_{p})=\hat{Y}_{p}.$ (54)
Taken together, $Y_{p}={i^{p}}_{\preceq p}^{\ !}(Q(e_{p}))={i^{p}}_{\preceq p}^{\ !}(\hat{Q}(e_{p}))={i^{p}}_{\preceq p}^{\ !}(\hat{Y}_{p})=\hat{Y}_{p}$. ∎
@@ 13.3 A pairing
Let $X=\mathbf{G}/\mathbf{B}$, $Y=\mathbf{G}/\mathbf{P}$ and let $\pi:X\to Y$ be the projection. It induces a map
$\pi^{*}:K_{\mathbf{T}}(Y)\to K_{\mathbf{T}}(X)$ and therefore, by subsection 4.3, a pairing
$K_{\mathbf{T}}(Y)\times K_{\mathbf{T}}(X)\to R(\mathbf{T}),\quad([{\mathcal{F}}],[{\mathcal{G}}])\mapsto\langle\pi^{*}[{\mathcal{F}}],[{\mathcal{G}}]\rangle.$
For $v,w\in W^{\mathbf{P}}$ with $vw_{0}\not\leq ww_{0}$ the pairing sends $([\mathcal{L}_{Y}(P(-e_{v}))],[\mathcal{O}_{{X}^{ww_{0}}}(-\partial{X}^{ww_{0}})])$ to $\beta_{v,ww_{0}}=0$. And it sends $([\mathcal{O}_{\overline{\mathbf{B}^{+}v\mathbf{P}}/\mathbf{P}}],[\mathcal{O}_{{X}^{ww_{0}}}(-\partial{X}^{ww_{0}})])$ to $\delta_{vw}$ for $v,w\in W^{\mathbf{P}}$. If ${\sf k}={\mathbb{C}}$ it follows that $[\mathcal{L}_{Y}(P(-e_{v}))]$ is in the $R(\mathbf{T})$-span of the Schubert classes $[\mathcal{O}_{\overline{\mathbf{B}^{+}w\mathbf{P}}/\mathbf{P}}]$ with $w\in W^{\mathbf{P}}$, $w\leq v$.
@@ 14.1 Borel-Weil-Bott theorem
    If $\lambda\in\overline{C}_{\mathbb{Z}}$ with $\lambda\notin X({\mathbf{T}})_{+}$, then
    ${\rm H}^{i}({\mathbf{G}}/{\mathbf{B}},\mathcal{L}_{w\cdot\lambda})=0$
    for all $w\in W$. Here $\cdot$ is the dot–action of the Weyl group $W$ on $X({\mathbf{T}})$.
@@
    If $\lambda\in\overline{C}_{\mathbb{Z}}\cap X({\mathbf{T}})_{+}$, then for all $w\in W$ and $i\in\mathbb{N}$
    ${\rm H}^{i}({\mathbf{G}}/{\mathbf{B}},\mathcal{L}_{w\cdot\lambda})={\rm H}^{0}({\mathbf{G}}/{\mathbf{B}},\mathcal{L}_{\lambda}),$
    if $i=l(w)$ and otherwise ${\rm H}^{i}({\mathbf{G}}/{\mathbf{B}},\mathcal{L}_{w\cdot\lambda})=0$. Here $l(w)$ is the length function.
@@ 14.2 Kempf vanishing theorem
[32, Part II, Proposition 4.5] and [21, Section 2.6] If $\lambda\in X({\mathbf{T}})_{+}$, then
${\rm H}^{i}({G}/{\mathbf{B}},\mathcal{L}_{\lambda})=0$
@@
By [5, Corollary 2.6, (ii)], ${\rm H}^{0}({\mathbb{G}}/{\mathbb{B}},{\mathbb{L}}_{\lambda})$ are free $\mathbb{Z}$-modules for $\lambda\in X({\mathbb{T}})_{+}$. Combined with the Universal coefficient Theorem 1.8 and Theorem 14.2, this gives ${\rm H}^{i}({\mathbb{G}}/{\mathbb{B}},{\mathbb{L}}_{\lambda})=0$ for $i>0$.
@@ 14.3 Kodaira vanishing theorem
[10, Theorem 1.2.9] Let $k$ be a field of characteristic $p>0$ and $X$ be a Frobenius split $k$-variety. Assume given an ample line bundle $\mathcal{L}$ on $X$. Then $H^{i}(X,\mathcal{L}\otimes\omega_{X})=0$ for $i>0$.
Generalized flag varieties $\mathbf{G}/\mathbf{P}$ are Frobenius split, [10, Theorem 2.2.5], thus Kodaira vanishing holds for these.
@@ 14.4 Demazure functors
By [32, Section 5.17-5.18], the Joseph-Demazure functors ${{\sf H}}_{w}$ from Section 2.2 can be related to functors which are defined geometrically. Let $R$ be the root system of a simple simply connected group $\mathbf{G}$ of rank $r$ and $\Pi=\{\alpha_{1},\dots,\alpha_{r}\}\subset R$ the simple roots. Let $s_{\alpha_{i}},i=1,\dots,r$ be the associated simple reflections of the Weyl group $W$. Given a simple root $\alpha$, let $\pi_{\alpha}:{\mathbf{G}}/{\mathbf{B}}\rightarrow{\mathbf{G}}/{\mathbf{P}}_{\alpha}$ denote the projection. Associated to $\alpha$ is an endofunctor ${\rm R}^{0}{\sf D}_{\alpha}:=\pi_{\alpha}^{\ast}{\rm R}^{0}{\pi_{\alpha}}_{\ast}$ of ${\mathsf{QCoh}}^{\mathbf{G}}({\mathbf{G}}/{\mathbf{B}})$ (the functor $\pi_{\alpha}^{\ast}$ is exact, the morphism $\pi_{\alpha}$ being smooth). Let $w$ be an element of the Weyl group. For a reduced expression of $w=s_{\alpha_{1}}s_{\alpha_{2}}\dots s_{\alpha_{n}}$ consider the functor ${\rm R}^{0}{\sf D}_{w}={\rm R}^{0}{\sf D}_{\alpha_{1}}\circ\dots\circ{\rm R}^{0}{\sf D}_{\alpha_{n}}=\pi_{\alpha_{1}}^{\ast}{\rm R}^{0}{\pi_{\alpha_{1}}}_{\ast}\dots\pi_{\alpha_{n}}^{\ast}{\rm R}^{0}{\pi_{\alpha_{n}}}_{\ast}$. The functor ${\rm R}^{0}{\sf D}_{w}$ is also a left exact functor.
Similarly to the case of ${{\sf H}}_{w}$, let ${\sf D}_{w}={\sf D}_{\alpha_{1}}\circ\dots\circ{\sf D}_{\alpha_{n}}=\pi_{\alpha_{1}}^{\ast}{\pi_{\alpha_{1}}}_{\ast}\dots\pi_{\alpha_{n}}^{\ast}{\pi_{\alpha_{n}}}_{\ast}$ denote the right derived functor of ${\rm R}^{0}{\sf D}_{w}$; it is an endofunctor of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}({\mathsf{QCoh}}^{\mathbf{G}}({\mathbf{G}}/{\mathbf{B}}))$ [^3] Note that the functor ${\sf D}_{w}$ is also defined on the derived category of non-equivariant coherent sheaves on $\mathbf{G}/\mathbf{B}$ and eventually will be considered on that category.. Since the morphisms $\pi_{\alpha_{i}}$’s are smooth and proper, the functors ${\sf D}_{w}$’s restrict to endofunctors of the bounded derived category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\sf Coh}^{\mathbf{G}}({\mathbf{G}}/{\mathbf{B}}))$ [^4] The same footnote as above applies..
Now set ${\mathbf{G}}:={\mathbf{B}},X={\mathbf{G}}$ (res., ${\mathbf{G}}:={\mathbf{P}}_{\alpha},X={\mathbf{G}}$) in the definition of the sheafification functor $\mathcal{L}$ and denote $\mathcal{L}_{\mathbf{B}}$ (resp., $\mathcal{L}_{{\mathbf{P}}_{\alpha}}$) the corresponding functors ${{\mathrm{Rep}}}({\mathbf{B}})\rightarrow{\mathsf{QCoh}}^{\mathbf{G}}({\mathbf{G}}/{\mathbf{B}})$ (resp., ${{\mathrm{Rep}}}({\mathbf{P}}_{\alpha})\rightarrow{\mathsf{QCoh}}^{\mathbf{G}}({\mathbf{G}}/{\mathbf{P}}_{\alpha}$)).
@@
Let $M\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}({{\mathrm{Rep}}}(\mathbf{B}))$ (resp., $M\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({{\mathrm{rep}}}(\mathbf{B}))$). Then there is an isomorphism $\mathcal{L}_{\mathbf{B}}({\rm R}{{\sf H}}_{w}(M))={\sf D}_{w}\mathcal{L}_{\mathbf{B}}(M)$.
@@
See [38, Lemma 5]. ∎
14.5 Demazure resolutions
Let $w\in W$ and $s_{\alpha_{i}}s_{\alpha_{i_{2}}}\dots s_{\alpha_{i_{k}}}$ be a reduced expression of $w$. Associated to $s_{\alpha_{i}}s_{\alpha_{i_{2}}}\dots s_{\alpha_{i_{k}}}$ is the Bott-Samelson-Demazure-Hansen resolution $Z_{w}$ of the Schubert variety $X_{w}$ associated to $w$, [10, Section 2.2] and [49, Section 1.2]. Both $X_{w}$ and $Z_{w}$ are acted upon by $\mathbf{B}$ and one can form $\mathbf{G}$-varieties by setting ${\mathcal{X}}_{w}=\mathbf{G}\times^{\mathbf{B}}X_{w}\subset X\times X$ and ${\mathcal{Z}}_{w}=\mathbf{G}\times^{\mathbf{B}}Z_{w}$. Then ${\mathcal{Z}}_{w}$ is a resolution of singularities of ${\mathcal{X}}_{w}$. Note that the variety ${\mathcal{Z}}_{w}$ depends on the choice of a reduced expression of $w$.
Theorem 14.6.
[49, Corollary 2.2.7]
${\sf D}_{w}=\Phi_{\mathcal{O}_{{\mathcal{X}}_{w}}}={q_{w}}_{\ast}p_{w}^{\ast}$ (55)
Proof.
To unburden the notation, let $X$ denote ${\mathbf{G}}/{\mathbf{B}}$. Occasionally though when there is a need, we will stress the notation ${\mathbf{G}}/{\mathbf{B}}$. For a simple root $\alpha_{i}$ let ${\mathbf{P}}_{\alpha_{i}}$ denote the minimal parabolic generated by ${\mathbf{B}}$ and $\mathbf{U}_{\alpha_{i}}$, and denote $Y_{i}:={\mathbf{G}}/{\mathbf{P}}_{\alpha_{i}}$ with $\pi_{\alpha_{i}}:X\rightarrow Y_{i}$. Given a simple root $\alpha_{i}$, we denote $\sf D_{i}$ the Demazure operator associated to $\alpha_{i}$, i.e. $\sf D_{i}:=\sf D_{\alpha_{i}}$. Throughout this section, we employ the notation of Section 1.5.2. The following theorem is the starting point for unraveling the category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({{\mathrm{rep}}}({\mathbf{B}}))$ as a module category over $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({{\mathrm{rep}}}(\mathbf{G}))$: